OpenStudy (anonymous):

does anyone knw answers for qns 3???

OpenStudy (anonymous):

FWCYDX

OpenStudy (anonymous):

means????

OpenStudy (anonymous):

thanks dude

OpenStudy (anonymous):

u knw answer for qns 5?formula pls

OpenStudy (anonymous):

for q5p1: il(t)=2*(1-e^(-t/tau) .... for il(0+) sub t=0 tau=(L/((R1+R2)||R3) .... for il(500m) sub t=500m tau= same as above...

OpenStudy (anonymous):

@kiraneswar93 , Q5 answers for ur parameters a) 0 b) 1.2642 c) 8 d) 8 e) 4.2821

OpenStudy (anonymous):

really thanks...so nice

OpenStudy (anonymous):

Click best response if answer is correct :)

OpenStudy (anonymous):

@kiraneswar93 Was the link helpful to u for Q2? U can give ur parameters, if it's same as mine, I can give solution directly.

OpenStudy (anonymous):

thanku the link was useful nd the parameters wer same for me